### v4

# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

Abstract
As global groundwater levels continue to decline rapidly, there is a growing need for advanced techniques to monitor and manage aquifers effectively. This study focuses on validating a numerical model using seismic data from a small-scale experimental setup designed to estimate water volume in a porous reservoir. Expanding on previous work with synthetic data, we analyze seismic data acquired from a controlled experimental site in Laukaa, Finland. By employing neural networks, we directly estimate water volume from seismic responses, bypassing the traditional need for separate determinations, for example, of reservoir water-table level and porosity. The study models wave propagation through a coupled poroviscoelastic-viscoelastic medium using a three-dimensional discontinuous Galerkin method. The proposed methodology is validated against experimental data, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
1 Introduction
Groundwater aquifers are facing unprecedented threats, with levels decreasing at alarming rates, often more than one meter per year in some areas—leading to significant long-term reductions. As a result, surface water flows that were previously sustained by groundwater are becoming seasonal or disappearing completely [10]. To ensure sustainable water extraction, it is crucial to have a better understanding of the location and extent of groundwater resources. Motivated by these challenges, this study explores the potential to estimate water volume in an artificial porous reservoir from seismic measurements using a combination of numerical modeling and neural networks, and validates the approach using controlled field data. Neural networks have been increasingly applied to enhance the analysis of seismic data, for a recent review, see [26]. By using neural network techniques, we aim to directly estimate water volume from seismic responses, avoiding the traditional need for separate determinations, for example, of reservoir water-table level and porosity.
Geophysical methods, including seismic techniques, are commonly used tools in the early stages of groundwater exploration and for ensuring sustainable extraction strategies [29, 12, 9]. They offer a cost-effective alternative to drilling, providing laterally continuous data across vast areas, by utilizing variations in material properties to detect subsurface features. Seismic methods, in particular, are well-suited for locating and monitoring water resources due to the higher seismic velocities exhibited by saturated materials compared to unsaturated ones [16]. The high resolution of seismic methods, both horizontally and vertically, enables detailed subsurface feature mapping.
Building upon previous works with synthetic data in both two (2D) [19] and three (3D) spatial dimensions [17], this study focuses on the estimation of water volume from seismic data collected at a controlled experimental site in Laukaa, Finland. The seismic data, obtained from an artificial porous sand pool using a drop-weight seismic source, were gathered during several acquisition campaigns following controlled changes in the water-table level. For the neural network-based water volume estimator, we first build a synthetic training database by simulating seismic wave propagation for different scenarios of the studied sand pool, including variations in physical parameters and water table level. The simulations employ a numerical wave propagation solver based on the methods presented in [7, 6, 17]. This synthetic data is then used to train neural networks, which are subsequently applied to the real seismic data to directly recover water volume, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
In our analysis of neural network-based estimates, we augment our approach by applying the Shapley additive explanation (SHAP) framework [22]. Employing the SHAP framework provides us with a deeper understanding of the estimation process at the receiver level. Specifically, we use SHAP analysis to evaluate the contribution of each receiver to the final water-volume estimate. The purpose is interpretability rather than receiver-array optimization.
In this study, we present a comprehensive framework for estimating water volume in a porous reservoir using seismic data and neural networks, validated through field experiments at the aforementioned Laukaa test site. Section 2 provides an overview of the site and details the seismic measurements conducted under varying water table conditions. In Section 3, we describe the construction of a 3D numerical model that replicates the test site, used for simulating synthetic seismograms based on Biot’s poroviscoelastic theory. Section 4 introduces the neural network architecture, including the data preprocessing steps, training methodology, and interpretability analysis using SHAP values. Results are shown in Section 5, with a focus on the water volume predictions and the SHAP analysis. In Section 6, we summarize our findings.
2 Site description and seismic measurements
The field measurements were conducted in a man-made sand pool at the Natural Resources Institute Finland (Luke) in Laukaa, Finland. Generally, the sand pool serves as a controlled environment to explore groundwater distribution with a knowledge of the media’s geometry and physical parameters. Although natural small-scale variability exists, the site can be reasonably approximated as a homogeneous and isotropic sand body for the purposes of our modelling. For conceptual clarity, we divide the domain into three approximate zones: an air-saturated porous region, a water-saturated porous region, and the surrounding elastic material. The pool is enclosed by an impermeable clay lining [28] and its dimensions are well defined, allowing controlled adjustments of the water-table level. This makes it possible to compute the true water volumes for each measurement configuration. The water-saturated region occupies the bottom of the pool (23.5 m by 8.2 m at 2 m depth), while the overlying air-saturated zone extends to approximately 29.5 m by 14.2 m.
The seismic measurements were conducted in June 2022. In our experiments, a metallic rod weight was dropped on a steel plate and served as a source of seismic waves, with an electric brake mechanism to control the rod’s release and prevent multiple hits. The experiments involved shots at 13 specific locations and were conducted at three different drop heights ($\rm{dh}$): 5, 10, and 15 cm. The different drop heights were used to vary the source amplitude and check for potential signal clipping at the nearest receivers.
For data acquisition, three-component (3C) 5 Hz geophones connected to 24-bit nodal seismic recorders were used, obtained from the Finnish Seismic Instrument Pool (FINNSIP) [14]. Various sources of ambient noise were present during the field campaign, including river currents, nearby construction activity, and occasional lawnmower operation. To reduce their influence, measurements were scheduled during rain-free and relatively quiet periods. The instruments were deployed along four receiver lines (14 receivers per line) with approximately 1.0 m inline and 1.7 m crossline spacing, plus one additional station placed at a random location, forming a fine-scale 3D acquisition layout. Exact receiver locations were surveyed using Differential Global Positioning System (DGPS) and incorporated directly into the numerical simulations. The seismic data were recorded at 4 kHz sampling rate.
Figure 1: Photo of experimental setup at Laukaa test site, showing the grid of geophones, weight-drop source, and wells to measure the water-table level.
Figure 1 shows the Laukaa test site, including the layout of the geophones, the weight-drop source, and the black pipes housing the water-table gauge wells. A total of 57 three-component geophones (blue) and their corresponding nodal recorders (white boxes) were deployed. To obtain ground-truth water-volume estimates, water-table levels were monitored using these gauge wells at several locations around the pool. The water level was adjusted by releasing water to the adjacent river, after which the system was allowed to stabilize before seismic acquisition. These auxiliary measurements were used exclusively for validating the estimated water volumes.
3 Modeling of the test site and field measurements
This section presents a three-dimensional (3D) synthetic model developed to simulate wave propagation and generate training data for the neural network. The model’s geometry for synthetic data generation replicates the Laukaa sand pool, capturing its structure and physical parameters. This replicated model serves as a tool for generating extensive synthetic training data for our neural network model. To generate synthetic seismograms, we employ Biot’s isotropic poroviscoelastic model to handle wave propagation in the porous medium. In a zone adjacent to the porous material, the isotropic viscoelastic model is utilized.
3.1 Governing equations
We model wave propagation in porous and elastic media using well-established partial differential equations (PDEs). Here, we summarize the main PDE systems and the associated physical parameters. More detailed derivations, numerical implementations, and attenuation models are available in [17] and references therein.
3.1.1 Biot’s poroelastic wave equation
The coupled dynamics of solid and fluid phases in porous media are described by Biot’s equations [3]
$\rho_{\mathrm{a}}\frac{\partial^{2}\mathbf{u}_{\rm s}}{\partial t^{2}}+\rho_{\mathrm{f}}\frac{\partial^{2}\mathbf{w}}{\partial t^{2}}$ $=\nabla\cdot\mathbf{T},$ (1)
$\rho_{\mathrm{f}}\frac{\partial^{2}\mathbf{u}_{\rm s}}{\partial t^{2}}+m\frac{\partial^{2}\mathbf{w}}{\partial t^{2}}+\frac{\eta}{k}\frac{\partial\mathbf{w}}{\partial t}$ $=\nabla\cdot\mathbf{T}_{\rm f},$ (2)
where $\mathbf{u}_{\rm s}$ is the solid displacement, and $\mathbf{w}=\phi(\mathbf{u}_{\rm f}-\mathbf{u}_{\rm s})$ is the relative fluid displacement (scaled by porosity $\phi$). Here, $\mathbf{u}_{\rm f}$ denotes the fluid displacement. The average density is $\rho_{\mathrm{a}}=(1-\phi)\rho_{\mathrm{s}}+\phi\rho_{\mathrm{f}}$, where $\rho_{\mathrm{f}}$ is the fluid density, $\rho_{\mathrm{s}}$ is the solid density, $\eta$ is the fluid viscosity, and $k$ is the permeability. The parameter $m=\rho_{\mathrm{f}}\tau/\phi$, with tortuosity $\tau$.
The total stress tensor $\mathbf{T}$ and fluid stress tensor $\mathbf{T}_{\rm f}$ are given by
$\mathbf{T}$ $=2\mu_{\mathrm{fr}}\mathbf{E}+\big(\lambda\operatorname{tr}(\mathbf{E})-\alpha_{\rm B}M\zeta\big)\mathbf{I},$ (3)
$\mathbf{T}_{\rm f}$ $=\big(\alpha_{\rm B}M\operatorname{tr}(\mathbf{E})-M\zeta\big)\mathbf{I},$ (4)
where $\mathbf{E}$ is the solid strain tensor and $\zeta=-\nabla\cdot\mathbf{w}$ is the fluid content variation. The parameter $\mu_{\mathrm{fr}}$ is the shear modulus of the frame, and $\mathbf{I}$ denotes the identity tensor.
The remaining parameters, $\lambda$, $\alpha_{\rm B}$, and $M$, are defined as
$\lambda$ $=\kappa_{\mathrm{fr}}+\alpha_{\rm B}^{2}M-\frac{2}{3}\mu_{\mathrm{fr}},$ (5)
$\alpha_{\rm B}$ $=1-\frac{\kappa_{\mathrm{fr}}}{\kappa_{\mathrm{s}}},$ (6)
$M$ $=\frac{\kappa_{\mathrm{s}}}{\alpha_{\rm B}-\phi(1-\kappa_{\mathrm{s}}/\kappa_{\mathrm{f}})},$ (7)
where $\kappa_{\mathrm{fr}}$, $\kappa_{\mathrm{s}}$, and $\kappa_{\mathrm{f}}$ are the frame, solid, and fluid bulk moduli, respectively.
3.1.2 Elastic wave equation
For purely elastic media, the displacement $\mathbf{u}_{\rm e}$ satisfies
$\rho_{\mathrm{e}}\frac{\partial^{2}\mathbf{u}_{\rm e}}{\partial t^{2}}=\nabla\cdot\mathbf{S},$ (8)
with stress tensor
$\mathbf{S}=2\mu_{\mathrm{e}}\mathbf{E}+\lambda_{\mathrm{e}}\operatorname{tr}(\mathbf{E})\mathbf{I},$ (9)
where $\rho_{\mathrm{e}}$ is the density, and $\mu_{\mathrm{e}}$, $\lambda_{\mathrm{e}}$ are the Lamé parameters. These parameters determine the compressional and shear wave speeds
$c_{P}=\sqrt{\frac{\lambda_{\mathrm{e}}+2\mu_{\mathrm{e}}}{\rho_{\mathrm{e}}}},\quad c_{S}=\sqrt{\frac{\mu_{\mathrm{e}}}{\rho_{\mathrm{e}}}}.$
At interfaces between elastic and poroelastic media, we follow the coupling strategy used in [7, 6]. The interface states are obtained by solving the corresponding elastic–poroelastic Riemann problem, subject to the physically required continuity conditions: continuity of solid velocity (normal and tangential) and continuity of total traction, while the normal fluid velocity vanishes on the elastic side.
3.1.3 Modeling attenuation
Attenuation in elastic and poroelastic media arises from different physical mechanisms. In both cases, stiffness attenuation is incorporated using a generalized Maxwell body (GMB) rheology [25]. The GMB introduces additional free parameters, specifically relaxation frequencies and memory variables, which govern the temporal behavior of stress relaxation and allow frequency-dependent attenuation. In elastic media, these parameters are tuned to match the desired quality factors $Q_{P}$ and $Q_{S}$. In poroelastic media, the quality factors $Q_{\mu_{\mathrm{fr}}}$, $Q_{\kappa_{\mathrm{fr}}}$, $Q_{\kappa_{\mathrm{s}}}$, and $Q_{\kappa_{\mathrm{f}}}$ are included in the model to account for frequency-dependent attenuation of the frame, solid, and fluid components. The full formulation of the GMB-based attenuation model, including how these memory variables are coupled with the governing PDEs and how the unrelaxed material parameters are derived, is provided in [17] and is not repeated here. For completeness, we note that the attenuation formulation used in [17] follows the poroelastic and viscoelastic GMB developments presented in [33, 32].
In poroelastic media, viscodynamic attenuation also plays a role, arising from viscous drag between fluid and solid phases. In this work, we operate within Biot’s low-frequency regime, where the dominant dissipation mechanism is frequency-independent fluid flow resistance. This results in a stiff PDE system due to the large contrast between relaxation and wave propagation timescales. We express the governing equations in first-order form using a velocity-strain formulation, leading to a hyperbolic system with stiff source terms. To handle this numerically, we apply Godunov splitting [21] to separate the dissipative and conservative components. Following [30], we solve the stiff dissipative terms analytically at each time step, while integrating the remaining hyperbolic system using the wave propagation software described in Section 3.4.
3.2 Synthetic model of the Laukaa test site
The applied synthetic model represents a finite section of the porous medium used for numerical simulations. It is a rectangular box with a length of 31.5 m, a width of 16.2 m, and a height of 2.75 m, with slightly rounded corners as shown in Fig. 2. It includes realistic air- and water-saturated subdomains, and a viscoelastic surrounding. The water-table level $z$ in the synthetic model is randomized as $z\sim\mathcal{U}(-120,-25)$ cm from the top surface.
Figure 2: A schematic of the problem geometry with the top showing cross-section and the bottom an oblique angle 3D view. Light-blue colour refers to air-saturated and dark-blue to water-saturated zone. Light-brown colour denotes the surrounding elastic material. The setup contains a total of 57 receivers, marked with red dots and one green dot, all of which are located on the ground surface. The green dot serves as the reference receiver and is located between the seventh receiver on lines 1 and 2. The source location is marked with a yellow star and placed close to the fourth receiver on line 1. This figure is motivated by the problem geometry presented in [17].
Our model includes a total of 57 receivers that record the solid velocity components in the horizontal ($y$) and vertical ($z$) directions, denoted by $v_{\rm s}$ and $w_{\rm s}$, respectively. These components represent the solid velocity in the model’s coordinate system (see Fig. 2). Notably, they do not correspond to isolated wave modes (e.g., P-, SV-, or SH-waves), but rather capture the full wavefield projected onto the model axes. As a result, the signals may contain contributions from multiple wave types, due to mode conversions and the complex nature of wave propagation in porous media. One additional receiver, marked in green, is located between the lines 1 and 2 which serves as a reference point. A more detailed discussion of this reference point is provided in Section 4.1.
Although the measurement data acquisition involved 13 shot locations, the water volume estimation is based on data from a single source location only, as used in the simulation study [17]. The location for the source was arbitrarily selected to be the closest to receiver 4 on Line 1. The seismic source is modeled as a vertical force pointing to the negative $z$-axis. Since the exact location of the source slightly varies between different measurements (distance from the closest geophone was measured with a ruler), the location was assumed to be uncertain in the numerical model when training the neural network. In practice, to account for uncertainty in repositioning the source between measurements at different water-table levels, we assume a 10 cm variability in the source location in both the $x$ and $y$ directions. This means that the source center location in the $(x,y)$-plane is randomized as $x\sim\mathcal{U}(4.4890,4.5890)$ m and $y\sim\mathcal{U}(-5.2142,-5.1142)$ m. Receivers and the source are placed on the ground surface.
3.3 Physical parameters
The computational domain consists of three layers: an elastic surrounding medium and two porous subdomains, which are water-saturated and air-saturated. The elastic medium is modeled as viscoelastic, with properties defined by solid density $\rho_{\mathrm{e}}$, pressure and shear wave speeds ($c_{P}$, $c_{S}$), and quality factors $Q_{P}$ and $Q_{S}$, which account for attenuation. The porous subdomains share the same solid matrix properties, including grain density, porosity, tortuosity, solid and frame moduli, and grain-size distribution. All of these material parameters are randomized from uniform distributions, with bounds provided in Table 1. The porous reservoir parameter values in Table 1 are based on site-specific information provided in the internal report (in Finnish and not publicly available) of the Geological Survey of Finland (GTK) as well as [28, 23]. The surrounding material (marked with light-brown in Fig. 2) is approximated by an idealized elastic medium whose parameters are selected based on the original structural descriptions of the site.
The distinction between the two porous layers lies in the fluid phase: the water-saturated domain uses fixed values for fluid density, bulk modulus, and viscosity specific to water, while the air-saturated domain uses different fixed values representative of air. These fluid parameters are also listed in Table 1.
Permeability $k$ is derived from the Darcy law
$\frac{k}{\eta}=\frac{K}{\rho_{f}g},$ (10)
where $g=9.81$ m s-2, and $K$ is the hydraulic conductivity estimated from grain-size parameters $d_{10}$ and $d_{50}$ following [2]
$K\,(\mbox{m/s})=a\left(I_{0}+0.025(d_{50}-d_{10})\right)^{2},$ (11)
with $a=1.505\times 10^{-3}$, and $I_{0}$ denoting the extrapolated lower-limit grain size [2]. The grain-size parameters are also randomized as shown in Table 1. We compute the permeability value for each material sample using the viscosity and density values for water. One must note, that the attenuation is modeled with three mechanisms in both porous reservoir and the surrounding medium, see [17].
Table 1: Material parameters used in the three-layer model. Layers: (1) elastic surrounding medium, (2) porous reservoir (water-saturated), (3) porous zone (air-saturated). Porous layers share the same solid matrix properties.
Layer Parameter Symbol (unit) Min Max
Layer 1: Elastic surrounding medium (viscoelastic)
Solid density $\rho_{\mathrm{e}}$ (kg m-3) 1400 1800
Pressure wave speed $c_{P}$ (m s-1) 1000 2000
Shear wave speed $c_{S}$ (m s-1) 400 800
Quality factor (P-wave) $Q_{P}$ 20 50
Quality factor (S-wave) $Q_{S}$ 20 50
Layers 2–3: Porous matrix (shared by both saturated zones)
Grain density $\rho_{\mathrm{s}}$ (kg m-3) 2400 2800
Solid bulk modulus $\kappa_{\mathrm{s}}$ (GPa) 45 55
Frame bulk modulus $\kappa_{\mathrm{fr}}$ (GPa) 0.008 0.05
Frame shear modulus $\mu_{\mathrm{fr}}$ (GPa) 0.002 0.04
Porosity $\phi$ (%) 30 40
Tortuosity $\tau$ 1.1 1.8
Quality factor (Shear modulus) $Q_{\mu_{\mathrm{fr}}}$ 15 50
Quality factor (Solid bulk modulus) $Q_{\kappa_{\mathrm{s}}}$ 80 120
Quality factor (Frame bulk modulus) $Q_{\kappa_{\mathrm{fr}}}$ 15 50
Quality factor (Fluid bulk modulus) $Q_{\kappa_{\mathrm{f}}}$ $\infty$
Grain size $d_{10}$ ($m$) 0.4 0.8
Grain size $d_{50}$ ($m$) 1.1 1.6
Layer 2 only: Fluid parameters (water-saturated)
Fluid density $\rho_{\mathrm{f}}$ (kg m-3) 1000
Fluid bulk modulus $\kappa_{\mathrm{f}}$ (GPa) 2.1025
[…]
Layer 3 only: Fluid parameters (air-saturated)
Fluid density $\rho_{\mathrm{f}}$ (kg m-3) 1.2
Fluid bulk modulus $\kappa_{\mathrm{f}}$ (MPa) 0.13628
[…]
The volume of stored water in the reservoir can be calculated by multiplying the volume of the water-saturated domain by the porosity. In this paper, the amount of water is calculated only from the water-saturated zone that is located exactly under the array of receivers. The current problem setup, with its simplified geometric structure, allows for the potential estimation of water volume across the entire reservoir. However, this cannot be assumed to be directly applicable to more complex and realistic aquifer models.
For the Laukaa test case studied in this paper, the parameter distributions shown in Table 1 are chosen to be representative of the geological properties observed in the region. However, it is important to note that while these ranges are designed to capture key variations, the neural network model may still perform well even if not all parameters are completely accounted for in the training data. This is because the network is able to learn relevant features from the available data, potentially allowing it to generalize and provide accurate estimates even in regions not explicitly covered by the training set. In this sense, the model can adapt to variations in parameters and still maintain robustness in its predictions [11].
3.4 Computation of seismic data
We utilize an in-house forward solver based on the discontinuous Galerkin (DG) method [13] and third-order Adams–Bashforth time-stepping [8] to compute synthetic seismograms. The solver is implemented in C/C++ and parallelized across CPU nodes using the Message Passing Interface (MPI). GPU acceleration is achieved through the OCCA library [24], which enables communication between the CPU and GPU as well as efficient execution on graphics hardware. The computational domain is discretized using tetrahedral elements, and accuracy is controlled via element size and the order of polynomial basis functions. All simulations were performed on the Puhti supercomputing environment maintained by CSC – IT Center for Science, Finland [5], using NVIDIA Volta V100 GPUs. Computational grids used in this work were built using COMSOL Multiphysics.
Two source time functions are used in our modeling: the first derivative of a Gaussian and the Ricker wavelet. Both wavelets are defined below using the same frequency and time delay parameters. We employ two source signatures for the following reason: the first derivative of a Gaussian wavelet is used to generate the training and validation datasets (Section 4.2), whereas the Ricker wavelet is used for the synthetic test set (Section 5.1). This intentional mismatch allows us to assess the ability of the predictive framework to generalize to an unseen source signature.
The first derivative of a Gaussian is defined as
$g=\frac{(t-t_{0})}{c}\exp\left(b\left((t-t_{0})^{2}-c^{2}\right)\right),$ (12)
where $b=-(f\pi)^{2}$ and $c=\sqrt{-0.5/b}$.
The Ricker wavelet is defined as
$g=\left(1+2b(t-t_{0})^{2}\right)\exp\left(b(t-t_{0})^{2}\right),$ (13)
with the same parameters $b$ and $t_{0}$, using a frequency $f=60$ Hz and delay $t_{0}=1.2/f$. Figure 3 (left) shows the source wavelet amplitudes as a function of time.
Figure 3: Left: The two source time functions used in the modeling, first derivative of a Gaussian (solid blue) and Ricker wavelet (dashed red). Right: Cross section of an example mesh used in the simulations.
To ensure accuracy across the simulations, we use polynomial basis functions of order four or five, depending on the dataset, and the mesh resolution is defined relative to the smallest expected wavelength in the model. Specifically, we use a minimum element density criterion of approximately 2.5 elements per shortest wavelength when generating test data (see Section 5.1), while the training and validation sets (see Section 4.2) use slightly lower resolutions, approximately 2.0 and 1.9 elements per wavelength, respectively. This ensures that the datasets have slightly different numerical accuracy, as intended. The purpose of using different discretizations, both in mesh resolution and basis order, is to avoid the so-called inverse crime: if the same forward model and numerical resolution were used for training, validation, and testing, the inversion could become artificially easier and unrealistically accurate. This principle is well established in the general theory of inverse problems [15]. The discretization differences introduce controlled numerical variability. These differences are not intended as noise but ensure that the test data are not generated with the exact same forward operator as the training data. An example mesh used in the synthetic test dataset simulations, shown in Fig. 3 (right), contains 158,180 tetrahedral elements and 28,407 vertices, illustrating the level of discretization applied.
4 Neural network-based characterization of water storage
4.1 Noise model, source function normalization, and data examples
Following [17] and denoting the measurement data vector and forward model as $\Upsilon=[v_{\rm s},w_{\rm s}]^{\top}$ and $\mathcal{A}$ respectively, the observation model is given by
$\Upsilon=\mathcal{A}(\bm{m})+e=X+e,$ (14)
where $\bm{m}$ contains all the physical and geometrical parameters of the applied model, such as porosity, tortuosity, and water-table level. The term $e$ represents additive noise components. The forward operator $\mathcal{A}$ is used to map the model parameters $\bm{m}$ to the synthetic seismic data vector $X$, simulated using the coupled viscoelastic-poroviscoelastic material model by the DG method in three spatial dimensions [6, 17].
To simulate measurement noise, we generate the noisy trace using the model
$X_{\textrm{noised}}=X+A|X|_{\max}\epsilon^{A}+B|X|\epsilon^{B},$ (15)
where $\epsilon^{A}$ and $\epsilon^{B}$ represent independent zero-mean Gaussian random variables, and $|X|_{\max}$ is the maximum absolute value of $X$. The two noise components correspond to additive white noise and amplitude-related noise, respectively, contributing to a diverse range of noise levels. This decomposition is motivated by practical considerations: in datasets with large dynamic range, a pure white-noise model disproportionately affects weaker signals while leaving stronger signals largely unperturbed. By including an amplitude-related term, the noise level scales with the signal, producing a more realistic distribution across all traces. While the additive component can be estimated from the data, the amplitude-dependent component cannot be robustly inferred. Therefore, parameters $A$ and $B$ are varied over wide intervals ($A\in[0.05,2]\%$, $B\in[0,2]\%$) to explore different noise conditions.
The seismic measurements were recorded continuously, but each source excitation produced a 3-second-long data segment, which was downloaded from the recorders based on microsecond-accurate, GPS-timed source initiations. For the neural network input, a 0.35-second window was extracted from both field and simulated data. This duration ensures that all significant wave reflections and late arrivals have attenuated to negligible amplitudes, preventing artificial truncation of energy and allowing each trace to end near zero amplitude.
After the waves produced by the seismic source have attenuated, we estimated the standard deviation of the Gaussian white noise component by analyzing a time window between 1 and 2 seconds. Comparing the estimated noise amplitude to peak signal amplitudes yielded an approximate noise level of $A_{\rm meas}\approx 0.4\%$.
In order to obtain a source-independent inversion, we use a deconvolution operation to remove the effect of the seismic source time function [20]. We transform the transient signals to the frequency domain and use data from a reference receiver (see Fig. 2) as the system response function. This allows us to rewrite the observation model (Eq. 14) as
$\frac{F_{\Upsilon}(\omega_{\ell})}{F_{\Upsilon_{\rm ref}}(\omega_{\ell})}$ $=$ $\frac{F_{X}(\omega_{\ell})}{F_{X_{{\rm ref}}}(\omega_{\ell})}+\hat{e}_{\ell},\qquad\ell=1,\ldots,N_{f},$ (16)
$\Leftrightarrow\quad\hat{V}$ $=$ $\hat{X}+\hat{e}$, (17)
where $F_{\Upsilon}$ is the Fourier transform of the measured seismic data at a given receiver, $F_{\Upsilon_{\rm ref}}$ is the Fourier transform of the data at the reference receiver, $F_{X}$ denotes the Fourier transform of the simulated (noise-free) model response, and $F_{X_{\rm ref}}$ is the corresponding simulated response at the reference receiver. Here, $\hat{e}$ is the noise term in the frequency-domain formulation after normalization, $\omega_{\ell}$ is the frequency at index $\ell$, and $N_{f}$ is the total number of frequencies. Note that the model (17) is applied separately to both velocity components, $v_{\rm s}$ and $w_{\rm s}$, at each receiver.
In practice, the division in Eq. (17) is implemented using the conjugate of $F_{X_{\rm ref}}(\omega_{\ell})$, leading to a denominator of the form $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}$. Since this denominator may approach zero, we apply a Wiener-type regularization. Following [31], the denominator $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}$ is replaced by $|F_{X_{\rm ref}}(\omega_{\ell})|^{2}+\epsilon\langle|F_{X_{\rm ref}}(\omega_{\ell})|^{2}\rangle,$ where $\epsilon=0.1$ is a stabilization constant. This regularization is applied directly within the frequency-domain ratio in Eq. (17). It does not replace Eq. (17) nor introduce any additional filtering procedure, it only serves to avoid numerical instabilities when the reference spectrum has small amplitudes.
To illustrate the behavior of real seismic data, Fig. 4 shows a sample shot gather from the field dataset, displaying the horizontal $v_{\rm s}$ and vertical $w_{\rm s}$ velocity components along with reference traces. The bottom row presents the real and imaginary parts of the Fourier-transformed and source-normalized records, which serve as inputs to the neural network. For this example, the water table was at 88.7-88.7 cm.
Figure 4: Measured seismic data: horizontal $v_{\rm s}$ component (left), vertical $w_{\rm s}$ component (right). Reference traces are shown in the top right panels.
Similarly, Fig. 5 presents an example of synthetic data with a water table at 72.10-72.10 cm and noise parameters $A=1\%$, $B=1\%$. Differences from real data arise from randomly selected material parameters and differing water-table levels. These examples illustrate the correspondence between measured and simulated records, and the type of inputs provided to the neural network.
Figure 5: Synthetic seismic data: horizontal $v_{\rm s}$ component (left), vertical $w_{\rm s}$ component (right). Reference traces for noisy and noiseless data are shown in the top panels.
4.2 Neural networks
We generated three distinct datasets for the neural network model: a training set consisting of 15,000 samples and a validation set of 3,000 samples. The training and validation datasets were created using fifth-order polynomials and the source wavelet (12) with the DG solver described in Section 3.4, which also details the mesh construction and numerical discretization. The values for the model input $\bm{m}$ were independently drawn from the probability distributions described in Sections 3.2 and 3.3 and used as inputs to the forward model $\mathcal{A}$ (see Eq. (17)). In addition to training and validation datasets, a separate test dataset comprising both synthetic and field measurement data was employed to evaluate the final performance of the neural network. This test dataset is described in detail in Section 5.1.
A fully connected neural network was employed in this study to determine the water volume using seismic data. By training on data with randomized model parameters $\bm{m}$ (see Eq. (14)), the network learns to focus on the relevant seismic signatures of water volume while marginalizing the influence of less significant factors. We down-sample the synthetic data to a sampling frequency of 4 kHz, then generate five copies of the clean training and validation datasets, each corrupted with Gaussian noise according to (15). This augmentation increases the size of both the training and validation datasets by a factor of five, which supports better generalization and robustness to measurement noise. The noise corrupted time domain data are transformed to the frequency domain according to model (17) and denoted as $\hat{X}_{\textrm{noised}}$ in the following. For each sample, we construct four input matrices of size $N_{r}\times N_{f}$, containing the real and imaginary parts of the two velocity components. Here, $N_{r}=56$ is the number of receivers and $N_{f}=21$ is the number of frequencies in the usable-energy band of the data (31.4–88.6 Hz with $\Delta f\approx 2.86$ Hz), see Fig. 5. These four matrices are stacked and reshaped into a single feature vector. This yields an input size of
$\text{frequency components}\times\text{number of receivers}\times\text{velocity components}\times 2=21\times 56\times 2\times 2=4704$
where the factor of 2 at the end accounts for the separate real and imaginary parts. Because the first layer is fully connected, each neuron receives all 4704 features, enabling the model to jointly learn from all four input channels. The network produces a single scalar output representing the estimated water volume.
The training process is implemented using the TensorFlow [1] and Keras [4] libraries. We employ the Adam optimizer [18] and utilize the RandomSearch algorithm from the Keras Tuner library [27] to optimize the model’s performance by exploring different hyperparameter configurations. The Keras Tuner is allowed to randomly select the activation function, learning rate, number of hidden layers, number of neurons per layer, and $L_{2}$ regularization penalty factor. The activation functions considered are “relu,” “sigmoid,” “tanh,” “selu,” “swish,” and “LeakyReLU,” while the number of hidden layers ranges from one to six. The number of neurons per layer varies between 100 and 5,000. The learning rate is selected from $1\times 10^{-3},1\times 10^{-4},1\times 10^{-5}$}, and the $L_{2}$ regularization penalty is chosen from $1\times 10^{-5},1\times 10^{-6},1\times 10^{-7},1\times 10^{-8}$}. For all networks in this study, the batch size is set to 256.
The training process is guided by minimizing the mean squared error (MSE) loss function, defined as
$\mathcal{L}(\theta)=\frac{1}{N_{{\rm train}}}\sum_{i=1}^{N_{{\rm train}}}\left(V_{{\rm true}}^{(i)}-\operatorname{NN}\left(\real{\hat{X}_{\textrm{noised}}^{(i)}},\mathbb{I}({\hat{X}_{\textrm{noised}}^{(i)}});\theta\right)\right)^{2}+\alpha\mathcal{R}(\theta),$ (18)
where $\operatorname{NN}$ denotes the neural network with weights and biases stored in $\theta$. Additionally, $N_{{\rm train}}$ represents the number of samples in the training dataset and $V_{{\rm true}}^{(i)}$ is the true water volume for the $i$-th sample. The term $\mathcal{R}(\theta)$ denotes the sum of the squared weights of the model and $\alpha$ is the $L_{2}$ regularization penalty coefficient. This loss function ensures that the predicted values closely approximate the true values. The model iteratively updates its weights and biases using the Adam optimizer to minimize this loss function. The convergence of the loss function over epochs is monitored to avoid overfitting. For a broader discussion on neural network training methodologies and optimization strategies, see [11, 34].
The best-performing network is selected based on MSE for validation. After testing various hyperparameter combinations, the final network consists of five hidden layers with 2,570 (layer 1), 3,920 (layer 2), 3,360 (layer 3), 2,730 (layer 4), and 3,400 (layer 5) neurons. It employs the LeakyReLU activation function, a learning rate of $1\times 10^{-5}$, and an $L_{2}$ regularization penalty factor of $1\times 10^{-5}$. The output layer uses a linear activation function, and early stopping is enabled during both hyperparameter tuning and training to enhance generalization and prevent overfitting.
The training process uses large, randomized datasets that broadly sample the parameter space, with noise added to help improve robustness. Hyperparameters are tuned via a random search, and early stopping is applied to reduce overfitting and support generalization to new data. Figure 6 illustrates the approach for estimating water volume from seismic data. The figure shows the sequential data processing steps, the neural network architecture with five hidden layers, and the training and validation loss curves that highlight the model’s learning and generalization over time.
Figure 6: Graphical illustration of the process for estimating water volume from seismic data. The left side depicts the pre-processing phase, including the deconvolution operation used to normalize the source wavelet effects. The right side presents the neural network-based estimation phase, showing the detailed architecture of the best-performing model. The network consists of five hidden layers with numbers of neurons $(m_{1},m_{2},m_{3},m_{4},m_{5})=(2570,3920,3360,2730,3400)$. Each hidden layer uses the LeakyReLU activation function, while the output layer employs a linear activation. The input size to the network is $n=4704$, corresponding to 2121 frequency components, 22 velocity components, 5656 receivers, and $2$ for the real and imaginary parts. The network produces a single scalar output estimating the water volume. The network was trained with a learning rate of $1\times 10^{-5}$ and an $L_{2}$ regularization penalty factor of $1\times 10^{-5}$. Early stopping was applied during both hyperparameter tuning and training to prevent overfitting and improve generalization. The top-right inset displays the loss curves over epochs for training and validation datasets, illustrating the model’s convergence behavior.
5 Results
The results are promising in terms of estimation accuracy. Notably, the estimation accuracy for the full receiver setup aligns closely with the true values and that of the supplementary synthetic database. However, it’s worth mentioning that one of the field data samples produced a significantly biased estimate when compared to the true value. Closer analysis of the data traces showed that the biased sample exhibited distinct differences, particularly in terms of the RMSE.
5.1 Predictions of water volume
We test the applicability of the trained neural network model via two different test datasets. These datasets are defined as:
  • Field Measurements: Field data were acquired at seven distinct water-table levels, ranging from a depth of –31.3 cm to –88.7 cm. At each level, three repeated measurements were recorded for each of three different drop heights, resulting in a dataset comprising $3\times 3\times 7=63$ observations. Using our knowledge of the actual water-table level and the geometry of the sand pool, we can calculate the volume of the water-saturated zone. Additionally, based on a previous study [28], we assume the nominal porosity of the material to be 35 per cent;
  • Synthetic Dataset: This test dataset contains a total of 3,000 samples, distinct from the training (15,000 samples) and validation (3,000 samples) datasets. The same prior distributions were used for material parameter sampling as in the training and validation datasets. Noise parameters $A$ and $B$ in (15) were fixed to 1% each, representing moderate noise levels. The DG simulations used fourth-order basis functions and the Ricker wavelet (see Eq. (13)) as the source function. As noted earlier, a modified mesh density and polynomial basis order were used to introduce controlled numerical variation from the training and validation datasets. Additionally, the Ricker wavelet is employed here, differing from the first derivative of a Gaussian wavelet used in the training phase, to test the neural network’s ability to generalize to unseen source signatures.
The Keras Tuner-optimized network architecture undergoes ten training runs, and the final water volume per sample result is determined as the average of these ten estimates. Figure 7 presents a comparison between the estimated (average) water volumes and their corresponding true values. The results corresponding to the 5 cm, 10 cm, and 15 cm drop heights are shown as three distinct groups, each containing the three repeated measurements $(\rm{dh}_{5}^{1,2,3})$, $(\rm{dh}_{10}^{1,2,3})$, and $(\rm{dh}_{15}^{1,2,3})$, respectively. These drop heights are colour coded in Figure 7, with red corresponding to 5 cm, green to 10 cm, and blue to 15 cm. To get a crude approximation for the uncertainty, we assumed that the porosity value in the field measurements database is uncertain in a sense that we assumed $\phi_{\rm{true}}\in[0.95,1.05]\times 35$ per cent, that can be used to compute the error bars shown for each estimate with real data. The figure shows also the estimates for the synthetic dataset. These results demonstrate the potential of using proposed neural network-based approach to recover the water volume.
Figure 7: Left: Estimated volumes of water as a function of the true value. Red denotes repeated measurements from drop height 5 cm ($\rm{dh}_{5}^{1,2,3}$), green from drop height 10 cm ($\rm{dh}_{10}^{1,2,3}$), and blue from drop height 15 cm ($\rm{dh}_{15}^{1,2,3}$). Right: Histogram of relative prediction errors, with the colour coding of the asterisks corresponding to the same drop heights as shown on the left.
These correlations are key to verifying the accuracy of our neural network model and demonstrate its potential for practical application. Quantitative comparison between seismic data-derived estimates and ground truth water volumes confirms the model’s reliability under controlled conditions. However, the occurrence of a significantly biased estimate for one field sample highlights the challenge posed by out-of-training-distribution data, likely caused by extreme or unexpected measurement noise, potentially hidden in the source-to-receiver offsets. This suggests the importance of sensitivity analysis and uncertainty quantification in further refining the method.
The results at water-table level -36.2 cm reveal a clear outlier in Fig. 7. To analyze the differences between repeated measurements, we utilize the RMSE. Let $d_{i}^{k}$ represent the normalized seismic data (see model (17)) for the $k$’th repeated measurement at drop heights $i\in\{5,10,15\}$ in the frequency domain, with the real and imaginary components stacked. The root mean square error (RMSE) error can now be expressed as follows:
${\rm RMSE}^{\ell-j}_{i}=\frac{\|d^{\ell}_{i}-d^{j}_{i}\|_{2}}{\sqrt{N}},$ (19)
where $\ell$ and $j$ are the indices for repeated measurements, and $N$ is the total number of values in the data vector. Table 2 lists the RMSE values between all possible combinations of input data. Specifically, the first drop height ($i=5$) exhibits significantly larger variations compared to the other two measurements. Additionally, the combination of drop iterations $\ell=2$ and $j=3$ yields comparable RMSE values for the measurements at $i=10$ and $i=15$.
Table 2: Root mean square errors (19) for different data combinations for measurements taken at a water-table level of -36.2 cm.
$i$ $\ell=1,j=2$ $\ell=1,j=3$ $\ell=2,j=3$
5 0.3780 0.3781 0.1399
10 0.1695 0.2112 0.1162
15 0.1401 0.1591 0.1070
5.2 SHAP analysis
We applied Shapley Additive Explanations (SHAP) analysis to the full receiver array neural network model to determine the significance of each receiver in estimating water volume. Determining Shapley values is an attribution problem, which means it involves determining the contribution of the prediction scores of a model for a specific sample input to its base features—in our case, the receivers. In simple terms, attribution to a base feature represents the importance of that feature to the prediction. For example, when attribution is applied to a model that estimates water volume, it helps us understand how influential each receiver is in determining the water volume.
10,000 randomly selected samples from the training dataset are used to train the deep explainer model for the SHAP software [22]. The explainer model is then applied to all samples in the field measurements database. After calculating the Shapley values, we compute the normalized mean absolute values for each receiver (see top panel of Fig. 8). The results for water volume indicate that the most contributing receivers are those closest to the seismic source. For comparison, we also applied the explainer model to 1,000 randomly selected samples from the synthetic database, revealing a similar distribution of the most contributing receivers.
Next, we constructed two new receiver configurations and trained the neural network model for the field measurements based on SHAP values. For the first configuration, we selected ten receivers having the largest SHAP values, and for the second configuration, we randomly selected ten receivers from the full sensor array (see bottom panel of Fig. 8). The KerasTuner optimized network with SHAP analysis-based receiver selection consists three hidden layers with 3,490 (layer 1), 3,780 (layer 2), and 2,580 (layer 3) neurons, a LeakyReLU activation function, a learning rate of 1e-4, and $L_{2}$ regularization penalty factor of 1e-6. Similarly, the randomly selected receiver selection lead to optimized network with two hidden layers with 3,490 (layer 1) and 2,580 (layer 2) neurons, a LeakyReLU activation function, a learning rate of 1e-4, and $L_{2}$ regularization penalty factor of 1e-6.
Figure 9 displays the estimated water volume as a function of the true water volume for both receiver configurations. The field measurement results reveal a significant impact on estimation accuracy when employing SHAP analysis-based receiver selection compared to random selection. However, with a synthetic database, this effect is not that significant.
Figure 8: Normalized average absolute Shapley values for predicting water volume from seismic data. The results for the test and real datasets are displayed. Higher Shapley values are correlated with receivers that are closer to the seismic source.
Figure 9: Left: Estimated volumes of water as a function of true value. Right: Histogram of relative prediction errors. Top panel shows the results for receiver configuration based on SHAP analysis and bottom panel for the randomly chosen receivers.
The normalized mean biases (NMB), mean absolute errors (MAE), and root mean square errors (RMSE) are used as evaluation metrics to quantitatively analyze the estimation results for all three receiver configurations, see Table 3. The table shows that the full receiver array results in the most accurate measures.
Table 3: The normalized mean biases (NMB), mean absolute errors (MAE), and root mean square errors (RMSE) were computed for both field measurement and synthetic test databases. It is important to note that, instead of averaging estimates from ten randomly initialized network trainings, they were treated as individual samples in this table. The first two rows show the results for the full receiver array, the following two lines for the ten receivers found most contributing according to SHAP analysis, and the last two rows the results for ten receivers selected randomly.
Receiver array setup Database NMB (%) MAE (m3) RMSE (m3)
full field meas. 1.2604 2.3259 2.9787
synthetic 1.5976 1.3801 1.7174
SHAP field meas. -1.3241 3.2449 4.1318
synthetic 0.5348 1.5146 1.8950
random field meas. -2.3320 6.5112 7.2178
synthetic 1.2344 1.5768 1.9630
The evaluation metrics indicate that the full receiver array provides the highest estimation accuracy, while SHAP value-based selection offers a practical compromise with somewhat reduced accuracy. Random receiver selection significantly degrades performance, underscoring the importance of receiver placement and data quality in seismic monitoring. These findings emphasize the value of explainability methods such as SHAP in guiding sensor array design and improving robustness of neural network-based estimations in real-world conditions.
6 Conclusions
In this study, we investigated the estimation of water volume in a reservoir using both field and synthetic seismic data. The reservoir’s controllable water-table level and known physical characteristics enabled direct validation of our neural network-based estimation approach. Synthetic training data were generated using a coupled poroviscoelastic–viscoelastic wave propagation model solved with a high-order discontinuous Galerkin method on a GPU cluster. The forward simulations used domain-wise homogeneous material properties and allowed for varying water-table levels within the geometry.
A fully connected neural network was trained to estimate water volume from frequency-domain seismic data normalized by a deconvolution method that accounts for unknown source wavelets. Training data were corrupted with noise levels matching field data conditions, ensuring robustness against realistic measurement noise.
The neural network results indicate that, with a full receiver array, water volume can potentially be estimated with good accuracy, approaching the known reservoir volumes. Analysis using SHAP values suggests that, in the present controlled setup, selecting the top ten most influential receivers can maintain performance comparable to the full array, while random receiver selection tends to degrade accuracy, particularly for field data. These findings highlight that receiver contributions influence estimation accuracy in this experiment. However, they are specific to the present setup and should not be assumed to generalize to other geological settings or problems.
Overall, this work presents a promising framework for seismic-based water volume estimation using neural networks. The findings emphasize the value of carefully chosen receiver configurations and suggest that models trained on synthetic data may generalize to controlled real-world conditions. While this study focuses on a simplified and well-constrained setup, the results point to the potential applicability of this approach in more complex environments in future work.
Acknowledgements
This work has been supported by the Research Council of Finland (the Finnish Centre of Excellence of Inverse Modeling and Imaging), Flagship of Advanced Mathematics for Sensing Imaging and Modelling (grant no. 358944), and the Research Council of Finland project 321761. The authors also wish to acknowledge the CSC – IT Center for Science, Finland, for generously sharing their computational resources. Special thanks to the Natural Resources Institute Finland (Luke) for sharing information about the sand pool and allowing us to carry out measurements on their premises. Authors would also like to thank Dr. Tuomo Savolainen from the Department of Technical Physics, University of Eastern Finland, for building the seismic source used in this work. The 3C nodal receivers used are a part of the Finnish national pool of seismic instruments [14].
Last, we would like to acknowledge our colleague, Kai Nyman, who contributed planning and implementation of the Laukaa measurements. Nyman passed away in 2022.
Appendix A Measured data, water-table level -36.2 cm
In this section, the data is shown for all measurements from drop height 5 cm at the water-table level -36.2 cm. Notably, the first measurement within this series led to biased water volume estimate, as elaborated in Section 5.1.
Figure 10: The figure depicts measured data traces, with the top panels displaying results in the time domain and the bottom panels showing the corresponding data converted into the frequency domain. These measurements were taken at a water-table level of -36.2 cm. Specifically, the figure presents the initial measurement out of a series of three, all conducted from a drop height of 5 cm.
Figure 11: Otherwise the same caption as in Fig. 10, but the data is for the second iteration of measurements.
Figure 12: Otherwise the same caption as in Fig. 10, but the data is for the third iteration of measurements.